k     3   3    
-  +  - = -
16    4   8


2p-56=6p
5b+22-4b=22+34b+7

Answer by mathdoc314(58)

This looks like two separate answers
k     3   3    
-  +  - = -
16    4   8

Multiply through by 16 which is the l.c.d. (least common denominator, or least common multiple of 16, 4, and 8).
16k/16 + 3*16/4 = 3*16/8
k + 3*4 = 3*2
k + 12 = 6
Subtract 12 from either side
k + 12 - 12 = 6 - 12
k = -6

Second problem: 2p-56=6p
I think you want to solve for p?
Subtract 2p from each side
-56 = 6p - 2p
-56 = 4p
Divide both sides by 4
-14 = p
That is the answer.
Third problem:
I think you want to solve for b?
5b+22-4b=22+34b+7
I will tell you the steps and you see if you can do them
Step 1. On the left side you can combine 5b - 4b = (5-4)b = b
Step 2. You can cancel 22 that is simply added on both sides of the equation
Step 3. That should leave you with b = 34b + 7
Step 4. Subtract 34b from both sides. 1b - 34b = -33 b
Step 5. Divide both sides by -33. This should give you b=-33/7